# Collaborative Web Navigation

### v2

@@ 1 Introduction
Agents supported by large language models (LLMs) have become increasingly capable of automating digital tasks such as web navigation (Zhou et al., 2023; Deng et al., 2024). While existing frameworks for web agents mostly focus on solo, autonomous agents Zheng et al. (2024b); Iong et al. (2024); Drouin et al. (2024), we argue that, for many practical tasks, users interact with the LLM agent for varied purposes such as supervision and collaboration, i.e., the copilot mode. While existing frameworks (Lù et al., 2024; Drouin et al., 2024; Wang et al., 2024a; Zheng et al., 2024b) mainly support users communicating with agents via natural language (NL) feedback, or recording actions of human users alone (Pan et al., 2024b), they do not support dynamic human-agent collaboration within a task session, where humans and LLM agents take actions alternately to solve tasks. We ask: How can we enable human-agent collaborative task-solving? and further, how do agents perform under such collaborative settings?
@@
Beyond agent web automation, CowPilot enables a wide range of use cases (§3), including web automation3.1), data collection for agent trajectories and user feedback (§3.2), as well as evaluations for single or multiple agents (§3.3).
@@ 2.1 The CowPilot System
Given an objective $o$ stated in natural language (NL) (e.g., book a flight from New York to Pittsburgh) for the web environment, we define two agents: one agent instantiated with an LLM policy $\pi_{\mathcal{L}}$, and one human agent $\pi_{\mathcal{H}}$. At each time step $t$, based on the observation $o_{t}$ from the environment state $s_{t}$, either the LLM agent or human agent generates an action $a_{t}$, formalized as $a_{t}=\pi(t,o_{t},a_{0:t-1})$. Executing $a_{t}$ on the environment results in a new state $s_{t+1}$ that gives observation $o_{t+1}$ that drives the next step. The two agents collectively generate a sequence of actions $a_{0:n}$ over $n$ steps, until it reaches a task termination condition, e.g., output STOP or a maximum number of steps. By default, the LLM agent starts generating actions $a_{0}^{\mathcal{L}}$ from $t=0$ unless intervened by the human agent $\mathcal{H}$. Table 1 shows the action space for LLM and human agents.
@@
Suggest-then-Execute under Human Supervision  At any time step, the human agent $\pi_{\mathcal{H}}$ can decide to take over by generating action $a_{t}^{\mathcal{H}}$. More concretely, the LLM agent $\pi_{\mathcal{L}}$ generates an action $a_{i}$ and presents it as a suggestion for the tentative next step to the user (Figure 2, 1), which includes a visual indicator highlighting the target element for the proposed action, accompanied by a textual explanation of the agent’s reasoning. This tentative step is presented to the human agent for at most five seconds, and is automatically executed if the human agent does not oppose. Otherwise, the human agent can choose to reject or pause the action (Figure 2, 3) and take over. They can also transfer the action back to the LLM-based agent by hitting the resume button (Figure 2, 4). This take-over-then-back process can be conducted unlimited times per task-solving session. This mechanism balances operational efficiency with user oversight, allowing users to intercept potential errors without the burden of manually approving every step.
Pause LLM Agent: Extract Human Actions  Whenever the human agent $\pi_{\mathcal{H}}$ rejects the LLM-proposed action, our CowPilot system starts tracking human activity on the websites, particularly what webpages and UI elements they interact with. To capture this micro-level metadata, we utilize HTML event listener [^1] https://developer.mozilla.org/en-US/docs/Web/API/EventTarget/addEventListener, that are attached to the interactive elements (e.g., text field, buttons, dropdown menu) in the current webpage and triggered each time the elements are accessed by the user.
Note that, actions captured by the HTML event listener can include noisy actions irrelevant to the task Siqueira and Baldochi (2018); Cheng and Kumar (2015), such as unintentional mouseover events. Hence, we transform the listener-captured actions to the LLM agent actions space, from Raw Human Action to Action column in Table 1, which are also coped with proper textual descriptions that can help the agent interpret user inputs effectively. To facilitate this transformation, we use an off-the-shelf LM (gpt-4o-2024-08-06) and provide the raw human actions as input. The model outputs the transformed and cleaned version of the actions. (Find the prompt for this transformation in §A.1.)
@@
Resume LLM Agent: Predict next Action using Human Input  If the human agent chooses to resume the LLM agent at any given step, our CowPilot stops tracking human actions and restarts LLM agent generation (Figure 2). Note that the LLM agent has access to all previous actions generated by itself and the human agent.
@@ 3 Use Cases of CowPilot
CowPilot has numerous potential use cases. We particularly highlight three use cases under the scope of this work.
@@ 3.1 Web Automation
CowPilot can be a standalone agent framework to automatically conduct web tasks for end users is implemented as a Chrome extension where all computations other than the LLM calls are handled locally with minimal storage requirement (<50MB). Any users can easily install CowPilot with just four clicks and use it with their personal API key. We use LiteLLM [^2] https://docs.litellm.ai/docs/ proxy server for our backend LLM, enabling CowPilot to support all models available via LiteLLM, including closed-source (e.g., GPT) and open-weight models (e.g., LLaMa). Depending on whether the user wants to participate in task-solving, our agent can operate in two modes: 1) Fully autonomous mode: the agent conducts a user-issued task start-to-end; 2) CoPilot mode: human and agent collaboratively solve a task, often useful for complex tasks where the agent is more prone to make mistakes.
@@ 3.2 Data Collection from Websites
We can track all actions conducted by LLM agents and humans. CowPilot can only track all actions from LLM and human agents, but also collect human feedback at both (i) step-level: whether the user judges the current step correctly leads to task success, and (2) task-level: whether the entire trajectory correctly solves the task. These rich data collections can easily facilitate various studies such as user behavior studies and advanced agent learning strategies.
@@ 3.3 Evaluation and Comparative Analysis of Agent Performance
CowPilot can be used to evaluate and compare agent performance. We support a wide range of open-weight and closed-source models served via LiteLLM. While this paper focuses on comparing GPT and LLaMa, the framework can easily extend to other open and closed-source models.
To evaluate a particular model, the user can select a model before initiating a task. Once the task is completed, CowPilot presents results evaluated in the metrics from §2.2. To compare different models on the same task, the user can re-do the task with different models, allowing for clear, unbiased comparisons under identical conditions.
@@ 4.2 Copilot Mode Requires Minimal Human Intervention
Despite the high task success rates, the GPT-based agent easily achieves the highest accuracy with an average of $1.1$ human steps, taking only $15.2\%$ of the entire trajectory. Instead, the LLM agent performs the majority, more precisely $84.8\%$ of task steps. Similarly, when shifting to the weaker LLaMa model, the human-llm collaboration process requires two times more human involvement, resulting in humans and LLM agents spending roughly similar amounts of effort, taking $4.47$ and $4.15$ respectively. Figure 3 shows the correlation between human step count and end-to-end task accuracy.
@@ 5.1 Web Agent Plugin
Chrome extensions, as adopted by tools like WebCanvas Pan et al. (2024b), WebOlympus Zheng et al. (2024b), OpenWebAgent Iong et al. (2024), and Taxy TaxyAI (2024), present a more user-friendly alternative. They are easy to install, lightweight, and integrate seamlessly into standard browsing environments, making them accessible to end-users. While similar to CowPilot, the extensions above lack features for fostering richer human-agent collaboration. Table 3 further compares how CowPilot with the existing frameworks by illustrating its novel features.
@@ 6 Limitation and Future Work
Currently CowPilot requires a human to act as an observer to oversee the task execution. This setup is intentional so that we can simulate task execution in live setting. We would like to extend our work so that it does not require constant human observation. Rather, we would detect the critical steps that require human observation only. In the future, we would extend CowPilot for a multi-LLM agent setup where we can simulate a user by a second LLM agent. Such setup would help us to approximate human decisions automatically using LLM autorater Pan et al. (2024a) and incorporate an active learning framework Bai et al. (2024). We acknowledge a potential ordering bias in the comparative evaluation of autonomous and CoPilot modes. We are currently conducting a large-scale study across a diverse demographic to assess and mitigate the impact of such biases.
@@
Web agents have significant potential to promote web accessibility and enhance user efficiency. However, their deployment raises important privacy and security concerns. For instance, tracking user actions may expose sensitive information, which could be exploited for malicious purposes (e.g. data theft). Additionally, agents may inadvertently perform harmful or irreversible actions (e.g. confirming financial transactions without explicit user consent). Beyond inadvertent risks, there is also the potential for intentional misuse, where malicious actors could exploit CowPilot for unethical purposes. We firmly discourage any such misuse of CowPilot. To balance accessibility with safety, we have chosen not to open-source our codebase. Instead, we will release the extension publicly through the Chrome Web Store. Users must provide their own API key for supported LLMs, or they can modify the extension to use a different LLM if they prefer, ensuring that their information is not shared with us. Future work can focus on addressing such safety risks and transparency, including developing robust safeguards to prevent unintended actions.

### v3

@@ 1 Introduction
Large language models (LLMs) have shown remarkable capabilities in automating digital tasks involving complex decision-making and information retrieval processes (Zhou et al., 2023; Deng et al., 2024). Despite these advancements, measuring the effectiveness of human-agent collaboration remains a challenging endeavor.
Previous studies have primarily focused on measuring task success and efficiency without adequately addressing the nuanced interactions between users and agents. In this follow-up work, we aim to address the following research questions:
  1. 1.
    How can we improve the evaluation of human-agent collaboration in web navigation tasks?
  2. 2.
    What factors contribute to successful or unsuccessful collaborations?
  3. 3.
    How do users’ interventions influence the overall effectiveness of the system?
While existing frameworks (Lù et al., 2024; Drouin et al., 2024; Wang et al., 2024a; Zheng et al., 2024b) mainly support users communicating with agents via natural language (NL) feedback, or recording actions of human users alone (Pan et al., 2024b), they do not support dynamic human-agent collaboration within a task session, where humans and LLM agents take actions alternately to solve tasks. We ask: How can we enable human-agent collaborative task-solving? and further, how do agents perform under such collaborative settings?
@@
Beyond agent web automation, CowPilot enables a wide range of use cases (§3), including web automation (§3.1), data collection for agent trajectories and user feedback (§3.2), as well as evaluations for single or multiple agents (§3.3).
@@ 2.1 The CowPilot System
Given an objective $o$ stated in natural language (NL) (e.g., book a flight from New York to Pittsburgh) for the web environment, we define two agents: one agent instantiated with an LLM policy $\pi_{\mathcal{L}}$, and one human agent $\pi_{\mathcal{H}}$. At each task step $t$, based on the observation $o_{t}$ from the environment state $s_{t}$, either the LLM agent or human agent generates an action $a_{t}$, formalized as $a_{t}=\pi(t,o_{t},a_{0:t-1})$. Executing $a_{t}$ on the environment results in a new state $s_{t+1}$ that gives observation $o_{t+1}$ that drives the next step. The two agents collectively generate a sequence of actions $a_{0:n}$ over $n$ steps, until it reaches a task termination condition, e.g., output STOP or a maximum number of steps. By default, the LLM agent starts generating actions $a_{0}^{\mathcal{L}}$ from $t=0$ unless intervened by the human agent $\pi_{\mathcal{H}}$. Table 1 shows the action space for LLM and human agents.
@@
Suggest-then-Execute under Human Supervision  At any time step, the human agent $\pi_{\mathcal{H}}$ can decide to take over by generating action $a_{t}^{\mathcal{H}}$. More concretely, the LLM agent $\pi_{\mathcal{L}}$ generates an action $a_{t}^{\mathcal{L}}$ and presents it as a suggestion for the tentative next step to the user (Figure 2, 1), which includes a visual indicator highlighting the target element for the proposed action, accompanied by a textual explanation of the agent’s reasoning. This tentative step is presented to the human agent for at most five seconds, and is automatically executed if the human agent does not oppose. Otherwise, the human agent can choose to reject or pause the action (Figure 2, 3) and take over. They can also transfer the action back to the LLM-based agent by hitting the resume button (Figure 2, 4). This take-over-then-back process can be conducted unlimited times per task-solving session. This mechanism balances operational efficiency with user oversight, allowing users to intercept potential errors without the burden of manually approving every step.
Pause LLM Agent: Extract Human Actions  Whenever the human agent $\pi_{\mathcal{H}}$ rejects the LLM-proposed action, our CowPilot system starts tracking human activity on the websites, particularly what webpages and UI elements they interact with. To capture this micro-level metadata, we utilize HTML event listener [^1] https://developer.mozilla.org/en-US/docs/Web/API/EventTarget/addEventListener, that are attached to the interactive elements (e.g., text field, buttons, dropdown menu) in the current webpage and triggered each time the elements are accessed by the user. To efficiently transfer user log data to the Chrome extension popup, we employ Chrome Message passing [^2] https://developer.chrome.com/docs/extensions/develop/concepts/messaging with a custom storage mechanism that is accessible by both the React AppState and chrome.localStorage [^3] https://developer.chrome.com/docs/extensions/reference/api/storage.
Note that, actions captured by the HTML event listener can include noisy actions irrelevant to the task Siqueira and Baldochi (2018); Cheng and Kumar (2015). Hence, we transform the listener-captured actions to the LLM agent actions space, from Raw Human Action to Action column in Table 1. Action also includes a textual description that helps the agent interpret user inputs effectively. To facilitate this transformation, we use an off-the-shelf LLM (gpt-4o-2024-08-06) and provide the raw human actions as input. The model outputs the transformed and cleaned version of the actions. (Find the prompt for this transformation in §A.1.)
@@
Resume LLM Agent: Predict next Action using Human Input  If the human agent chooses to resume the LLM agent at any given step, CowPilot stops tracking human actions and restarts LLM agent generation (Figure 2). Note that the LLM agent has access to all previous actions generated by itself and the human agent.
@@ 3 Use Cases of CowPilot
CowPilot can serve as a useful tool for agentic research on understanding and annotating human-agent collaborative behavior. We particularly highlight three use cases in the current scope of this work.
@@ 3.1 Web Automation
CowPilot can be a standalone agent framework to automatically conduct web tasks for end users. CowPilot is implemented as a Chrome extension where all computations other than the LLM calls are handled locally with minimal storage requirement (<50MB). Any users can easily install CowPilot with just four clicks and use it using their personal API key. We use LiteLLM [^4] https://docs.litellm.ai/docs/ proxy server for our backend LLM, enabling CowPilot to support all models available via LiteLLM, including closed-source (e.g., GPT) and open-weight models (e.g., LLaMa). Depending on whether the user wants to participate in task-solving, our agent can operate in two modes: 1) Fully autonomous mode: the agent conducts a user-issued task start-to-end; 2) CoPilot mode: human and agent collaboratively solve a task, often useful for complex tasks where the agent is more prone to make mistakes.
@@ 3.2 Data Collection from Websites
CowPilot can track all actions conducted by LLM agents and humans. Besides, it can also collect human feedback at both (i) step-level: whether the user judges that the current step to lead to task success, and (2) task-level: whether the entire trajectory correctly solves the task. These rich data collections can easily facilitate various studies such as user behavior studies and advanced agent learning strategies.
@@ 3.3 Evaluation and Comparative Analysis of Agent Performance
CowPilot can be used to evaluate and compare agent performance. We support a wide range of open-weight and closed-source models served via LiteLLM. While this paper focuses on comparing GPT and LLaMa, the framework can easily extend to other open and closed-source models accessible through LiteLLM.
To evaluate a particular model, the user can select a model before initiating a task. Once the task is completed, CowPilot presents the task summary containing the results on our proposed evaluation metrics from §2.2. To compare different models on the same task, the user can re-do the task with different models, allowing for clear, unbiased comparisons under identical conditions. We purposefully prevent the users from changing the LLM backbone when a task is running, so that the results are not biased by the other LLMs in competition.
@@ 4.2 Copilot Mode Requires Minimal Human Intervention
Despite the high task success rates, the GPT-based agent easily achieves the highest accuracy with an average of $1.1$ human steps, taking only $15.2\%$ of the entire trajectory. Instead, the LLM agent performs the majority, more precisely $84.8\%$ of task steps. Similarly, when shifting to the weaker LLaMa model, the human-llm collaboration process requires two times more human involvement, resulting in humans and LLM agents spending roughly similar amounts of effort, taking $4.77$ and $4.15$ respectively. Figure 3 shows the correlation between human step count and end-to-end task accuracy.
@@ 5.1 Web Agent Plugin
Chrome extensions, as adopted by tools like WebCanvas Pan et al. (2024b), WebOlympus Zheng et al. (2024b), OpenWebAgent Iong et al. (2024), and Taxy TaxyAI (2024), present a more user-friendly alternative. They are easy to install, lightweight, and integrate seamlessly into standard browsing environments, making them accessible to end-users. While similar to CowPilot, the extensions above lack features for fostering richer human-agent collaboration. Table 3 further compares how CowPilot with the existing frameworks by illustrating its novel features. The comparison is based their affordance in six categories: 1) Live, self-hosted website (e.g., WebArena); 2) Dynamic website that changes over time (e.g., shopping website); 3) Collection of end-to-end (i.e., human-only mode) annotation; 4) Interaction between human and agent; 5) Co-execution capability of human and agent when a task is ongoing; and 6) Interactive Evaluation of Agent performance.
@@ 6 Limitation and Future Work
Currently CowPilot requires a human to act as an observer to oversee the task execution. This setup is intentional so that we can simulate task execution in live setting. We would like to extend our work so that it does not require constant human observation. Rather, we would detect the critical steps that require human observation only. In the future, we would extend CowPilot for a multi-LLM agent setup where we can simulate a user by a second LLM agent. Such setup would help us to approximate human decisions automatically using LLM autorater Pan et al. (2024a) and incorporate an active learning framework Bai et al. (2024). We acknowledge a potential ordering bias in the comparative evaluation of autonomous and CoPilot modes. Moreover, due to the human-centric nature of CowPilot’s design, the results and intervention pattern might vary based on the user. We are currently conducting a large-scale study across a diverse demographic and diverse websites to assess and mitigate the impact of such behavioral patterns.
@@
Web agents have significant potential to promote web accessibility and enhance user efficiency. However, their deployment raises important privacy and security concerns. For instance, tracking user actions may expose sensitive information, which could be exploited for malicious purposes (e.g. data theft). Additionally, agents may inadvertently perform harmful or irreversible actions (e.g. confirming financial transactions without explicit user consent). Beyond inadvertent risks, there is also the potential for intentional misuse, where malicious actors could exploit CowPilot for unethical purposes. We firmly discourage any such misuse of CowPilot. To ensure safety, we have chosen not to open-source our codebase until we perform an extensive user study on CowPilot. For now, we are only releasing the final chrome extension package. Users must provide their own API key for supported LLMs, or they can modify the extension to use a different LLM if they prefer, ensuring that their information is not shared with us. Future work can focus on addressing such safety risks and transparency, including developing robust safeguards to prevent unintended actions.